# Stochastic Encodings for Active Feature Acquisition

### v1

@@ 4 Limitations of CMI Maximization
[…]
@@ 5.1 Architecture and Training
Training.   The SEFA architecture is trained in a supervised manner, minimizing the predictive negative log-likelihood, avoiding potential issues associated with training using RL. We also impose an information bottleneck regularization term on the latent space $I(X_{S};Z)$ (Tishby et al., 1999), minimizing how much information about the features is encoded in the latent variable. The regularization term is implemented using a variational upper bound (Alemi et al., 2017), as the expected KL divergence between $p_{\theta}(\mathbf{z}|\mathbf{x}_{S})$ and a prior $p(\mathbf{z})$. We use $\mathcal{N}(0,1)$ as the prior so that the KL divergence has a closed-form expression. This is multiplied by $\beta$ to balance predictive performance with regularization. Additionally, to train SEFA to make predictions on arbitrary feature subsets, for a given batch, we first subsample the feature subset, such that each available feature has probability 0.50.50.5 of being removed. This gives the loss
$L=\mathop{{}{\mathbb{E}}}_{p_{\text{Data}}(\mathbf{x}_{S},y)}$ $\mathop{{}{\mathbb{E}}}_{p_{\text{Bernoulli}}(S^{\prime})}\bigg{[}-\log\big{(}p_{\theta,\phi}(y|\mathbf{x}_{S\cap S^{\prime}})\big{)}$ (1)
@@ 5.2 Acquisition
$r(c,\mathbf{z},i)=\frac{||\mathbf{g}_{\mathcal{G}_{i}}||_{2}}{\sum_{j}||\mathbf{g}_{\mathcal{G}_{j}}||_{2}},~~~~~\mathbf{g}=\nabla_{\mathbf{z}}p_{\phi}(Y=c|\mathbf{z}).$ (3)
@@
In order to convert the gradients of $\mathbf{z}$ to feature scores, we calculate the length of the gradient in a feature’s associated latent components, $||\mathbf{g}_{\mathcal{G}_{i}}||_{2}$. Note that this is only possible because we used feature-wise encoders rather than one fully connected encoder. Finally, we normalize scores to treat each latent sample equally, removing the effect of the total gradient length. See Figure 1 for an example calculation.
@@ 5.3 Understanding the Latent Space
Revisiting the Loss.   The predictive part of the loss in (1) is given by the negative log-likelihood $-\log\mathop{{}{\mathbb{E}}}_{p_{\theta}(\mathbf{z}|\mathbf{x}_{S})}p_{\phi}(y|\mathbf{z})$. It is standard to take one latent sample to calculate this. However, we take multiple samples for two reasons. (1) The predictor network must see enough latent samples during training to learn to adapt based on different samples during acquisition. (2) If only one sample is taken for a given $\mathbf{x}_{S}$, the predictive loss encourages each sample to output the same prediction. Either the encoder $p_{\theta}(\mathbf{z}|\mathbf{x}_{S})$ reduces the latent standard deviation, or the predictor $p_{\phi}(y|\mathbf{z})$ predicts the same label distribution for the different latent samples. Either way, there is a lack of diversity in the latent space predictions. If multiple samples are taken, this allows individual samples’ predictions to vary more. This will not improve the predictions, but improves the acquisition since a more diverse set of possible realizations are used in the inner loop of the objective. We verify this empirically in Section 6 and Appendix C.
@@ 6 Experiments
ACFlow 7.730±0.1397.730$\pm$ 0.1397.730 ± 0.139 7.527±0.2547.527$\pm$ 0.2547.527 ± 0.254 9.194±0.2789.194$\pm$ 0.2789.194 ± 0.278
DIME 4.079±0.0644.079$\pm$ 0.0644.079 ± 0.064 4.581±0.2174.581$\pm$ 0.2174.581 ± 0.217 5.667±0.0385.667$\pm$ 0.0385.667 ± 0.038
EDDI 9.197±0.2039.197$\pm$ 0.2039.197 ± 0.203 9.214±0.4159.214$\pm$ 0.4159.214 ± 0.415 9.794±0.1869.794$\pm$ 0.1869.794 ± 0.186
Fixed MLP 6.009±0.0006.009$\pm$ 0.0006.009 ± 0.000 5.996±0.0005.996$\pm$ 0.0005.996 ± 0.000 7.999±0.0007.999$\pm$ 0.0007.999 ± 0.000
GDFS 4.568±0.2194.568$\pm$ 0.2194.568 ± 0.219 4.484±0.1594.484$\pm$ 0.1594.484 ± 0.159 5.587±0.2015.587$\pm$ 0.2015.587 ± 0.201
GSMRL 5.570±0.1275.570$\pm$ 0.1275.570 ± 0.127 6.227±0.1856.227$\pm$ 0.1856.227 ± 0.185 8.199±0.0678.199$\pm$ 0.0678.199 ± 0.067
Opportunistic RL 4.201±0.0414.201$\pm$ 0.0414.201 ± 0.041 4.846±0.0214.846$\pm$ 0.0214.846 ± 0.021 5.850±0.0725.850$\pm$ 0.0725.850 ± 0.072
Random 9.484±0.0069.484$\pm$ 0.0069.484 ± 0.006 9.499±0.0059.499$\pm$ 0.0059.499 ± 0.005 9.987±0.0089.987$\pm$ 0.0089.987 ± 0.008
VAE 6.589±0.0886.589$\pm$ 0.0886.589 ± 0.088 6.667±0.1406.667$\pm$ 0.1406.667 ± 0.140 7.888±0.0657.888$\pm$ 0.0657.888 ± 0.065
@@
$\beta=0$ 4.520±0.0824.520$\pm$ 0.0824.520 ± 0.082 4.578±0.1094.578$\pm$ 0.1094.578 ± 0.109 5.716±0.1045.716$\pm$ 0.1045.716 ± 0.104
1 Acquisition Sample 4.683±0.0304.683$\pm$ 0.0304.683 ± 0.030 4.862±0.0354.862$\pm$ 0.0354.862 ± 0.035 5.700±0.0275.700$\pm$ $0$ 0.027
1 Train Sample 4.421±0.1744.421$\pm$ 0.1744.421 ± 0.174 4.713±0.1604.713$\pm$ 0.1604.713 ± 0.160 5.188±0.1075.188$\pm$ 0.1075.188 ± 0.107
Deterministic Encoder 4.593±0.2364.593$\pm$ 0.2364.593 ± 0.236 4.773±0.2284.773$\pm$ 0.2284.773 ± 0.228 5.744±0.0345.744$\pm$ 0.0345.744 ± 0.034
Feature Space Calculation 5.111±0.0705.111$\pm$ 0.0705.111 ± 0.070 5.461±0.1285.461$\pm$ 0.1285.461 ± 0.128 5.977±0.0525.977$\pm$ 0.0525.977 ± 0.052
WO Normalization 4.036±0.0094.036$\pm$ 0.0094.036 ± 0.009 4.104±0.0084.104$\pm$ 0.0084.104 ± 0.008 5.101±0.0155.101$\pm$ 0.0155.101 ± 0.015
@@ 6.2 Datasets with Unknown Feature Orderings
ACFlow 0.899±0.0010.899$\pm$ 0.0010.899 ± 0.001 0.823±0.0130.823$\pm$ 0.0130.823 ± 0.013 0.558±0.0100.558$\pm$ 0.0100.558 ± 0.010 0.881±0.0180.881$\pm$ 0.0180.881 ± 0.018
DIME 0.901±0.0010.901$\pm$ 0.0010.901 ± 0.001 0.907±0.0020.907$\pm$ 0.0020.907 ± 0.002 0.661±0.0020.661$\pm$ 0.0020.661 ± 0.002 0.951±0.0010.951$\pm$ 0.0010.951 ± 0.001
EDDI 0.764±0.0050.764$\pm$ 0.0050.764 ± 0.005 0.705±0.0100.705$\pm$ 0.0100.705 ± 0.010 0.412±0.0130.412$\pm$ 0.0130.412 ± 0.013 0.842±0.0090.842$\pm$ 0.0090.842 ± 0.009
Fixed MLP 0.883±0.0010.883$\pm$ 0.0010.883 ± 0.001 0.909±0.0010.909$\pm$ 0.0010.909 ± 0.001 0.658±0.0020.658$\pm$ 0.0020.658 ± 0.002 0.954±0.0000.954$\pm$ 0.0000.954 ± 0.000
GDFS 0.900±0.0000.900$\pm$ $0$ 0.000 0.907±0.0010.907$\pm$ 0.0010.907 ± 0.001 0.653±0.0020.653$\pm$ 0.0020.653 ± 0.002 0.949±0.0000.949$\pm$ 0.0000.949 ± 0.000
GSMRL 0.891±0.0010.891$\pm$ 0.0010.891 ± 0.001 0.879±0.0060.879$\pm$ 0.0060.879 ± 0.006 0.638±0.0030.638$\pm$ 0.0030.638 ± 0.003 0.946±0.0010.946$\pm$ 0.0010.946 ± 0.001
Opportunistic RL 0.901±0.0000.901$\pm$ 0.0000.901 ± 0.000 0.910±0.0000.910$\pm$ 0.0000.910 ± 0.000 0.657±0.0010.657$\pm$ 0.0010.657 ± 0.001 0.953±0.0000.953$\pm$ 0.0000.953 ± 0.000
Random 0.699±0.0010.699$\pm$ 0.0010.699 ± 0.001 0.816±0.0030.816$\pm$ 0.0030.816 ± 0.003 0.569±0.0030.569$\pm$ 0.0030.569 ± 0.003 0.912±0.0010.912$\pm$ 0.0010.912 ± 0.001
VAE 0.901±0.0010.901$\pm$ 0.0010.901 ± 0.001 0.878±0.0020.878$\pm$ 0.0020.878 ± 0.002 0.633±0.0050.633$\pm$ 0.0050.633 ± 0.005 0.925±0.0020.925$\pm$ 0.0020.925 ± 0.002
@@
ACFlow 0.667±0.0030.667$\pm$ 0.0030.667 ± 0.003 0.652±0.0020.652$\pm$ 0.0020.652 ± 0.002 0.542±0.0060.542$\pm$ 0.0060.542 ± 0.006 0.711±0.0090.711$\pm$ 0.0090.711 ± 0.009
DIME 0.731±0.0020.731$\pm$ 0.0020.731 ± 0.002 0.703±0.0020.703$\pm$ 0.0020.703 ± 0.002 0.670±0.0070.670$\pm$ 0.0070.670 ± 0.007 0.805±0.0020.805$\pm$ 0.0020.805 ± 0.002
EDDI 0.572±0.0030.572$\pm$ 0.0030.572 ± 0.003 0.604±0.0010.604$\pm$ 0.0010.604 ± 0.001 0.563±0.0110.563$\pm$ 0.0110.563 ± 0.011 0.634±0.0050.634$\pm$ 0.0050.634 ± 0.005
Fixed MLP 0.708±0.0010.708$\pm$ 0.0010.708 ± 0.001 0.690±0.0010.690$\pm$ 0.0010.690 ± 0.001 0.685±0.0030.685$\pm$ 0.0030.685 ± 0.003 0.799±0.0040.799$\pm$ 0.0040.799 ± 0.004
GDFS 0.732±0.0010.732$\pm$ 0.0010.732 ± 0.001 0.692±0.0020.692$\pm$ 0.0020.692 ± 0.002 0.671±0.0050.671$\pm$ 0.0050.671 ± 0.005 0.797±0.0020.797$\pm$ 0.0020.797 ± 0.002
GSMRL 0.701±0.0020.701$\pm$ 0.0020.701 ± 0.002 0.683±0.0010.683$\pm$ 0.0010.683 ± 0.001 0.665±0.0020.665$\pm$ 0.0020.665 ± 0.002 0.781±0.0030.781$\pm$ 0.0030.781 ± 0.003
Opportunistic RL 0.740±0.0000.740$\pm$ 0.0000.740 ± 0.000 0.708±0.0000.708$\pm$ 0.0000.708 ± 0.000 0.706±0.0040.706$\pm$ 0.0040.706 ± 0.004 0.838±0.0020.838$\pm$ 0.0020.838 ± 0.002
Random 0.661±0.0010.661$\pm$ 0.0010.661 ± 0.001 0.648±0.0010.648$\pm$ 0.0010.648 ± 0.001 0.647±0.0050.647$\pm$ 0.0050.647 ± 0.005 0.753±0.0030.753$\pm$ 0.0030.753 ± 0.003
VAE 0.716±0.0010.716$\pm$ 0.0010.716 ± 0.001 0.685±0.0010.685$\pm$ 0.0010.685 ± 0.001 0.690±0.0040.690$\pm$ 0.0040.690 ± 0.004 0.800±0.0030.800$\pm$ $0$ 0.003
@@ Acknowledgements
The results in the TCGA experiment are based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga. We thank the ICML reviewers for their comments and suggestions. We also thank Lucie Charlotte Magister and Iulia Duta for providing feedback. Alexander Norcliffe is supported by a GSK plc grant. Changhee Lee is supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00358602) and the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2019-II190079), Artificial Intelligence Graduate School Program (Korea University).
@@ Appendix C Real Data Ablations
$\beta=0$ $\underline{0.906\pm 0.000}$ 0.917±0.0010.917$\pm$ 0.0010.917 ± 0.001 $\underline{0.680\pm 0.005}$ 0.957±0.0000.957$\pm$ 0.0000.957 ± 0.000
1 Acquisition Sample 0.889±0.0010.889$\pm$ 0.0010.889 ± 0.001 0.912±0.0010.912$\pm$ 0.0010.912 ± 0.001 0.667±0.0040.667$\pm$ 0.0040.667 ± 0.004 0.952±0.0000.952$\pm$ 0.0000.952 ± 0.000
1 Train Sample 0.901±0.0020.901$\pm$ 0.0020.901 ± 0.002 0.912±0.0030.912$\pm$ 0.0030.912 ± 0.003 0.676±0.0060.676$\pm$ 0.0060.676 ± 0.006 0.953±0.0000.953$\pm$ 0.0000.953 ± 0.000
Deterministic Encoder 0.900±0.0000.900$\pm$ $0$ 0.000 0.893±0.0060.893$\pm$ 0.0060.893 ± 0.006 0.667±0.0050.667$\pm$ 0.0050.667 ± 0.005 0.944±0.0020.944$\pm$ 0.0020.944 ± 0.002
Feature Space Calculation 0.897±0.0020.897$\pm$ 0.0020.897 ± 0.002 0.880±0.0030.880$\pm$ 0.0030.880 ± 0.003 0.600±0.0040.600$\pm$ $0$ 0.004 0.914±0.0040.914$\pm$ 0.0040.914 ± 0.004
WO Copula $\underline{0.905\pm 0.000}$ 0.917±0.0020.917$\pm$ 0.0020.917 ± 0.002 0.675±0.0020.675$\pm$ 0.0020.675 ± 0.002 0.952±0.0010.952$\pm$ 0.0010.952 ± 0.001
WO Normalization $\underline{0.905\pm 0.000}$ 0.919±0.0010.919$\pm$ 0.0010.919 ± 0.001 0.676±0.0050.676$\pm$ 0.0050.676 ± 0.005 0.957±0.0000.957$\pm$ 0.0000.957 ± 0.000
WO Prob Weighting 0.904±0.0010.904$\pm$ 0.0010.904 ± 0.001 0.919±0.0010.919$\pm$ 0.0010.919 ± 0.001 0.674±0.0050.674$\pm$ 0.0050.674 ± 0.005 0.957±0.0000.957$\pm$ 0.0000.957 ± 0.000
SEFA (full) 0.904±0.0010.904$\pm$ 0.0010.904 ± 0.001 0.919±0.0010.919$\pm$ 0.0010.919 ± 0.001 0.676±0.0050.676$\pm$ 0.0050.676 ± 0.005 0.957±0.0000.957$\pm$ 0.0000.957 ± 0.000
@@
$\beta=0$ 0.759±0.0000.759$\pm$ 0.0000.759 ± 0.000 0.719±0.0000.719$\pm$ 0.0000.719 ± 0.000 0.706±0.0030.706$\pm$ 0.0030.706 ± 0.003 0.842±0.0020.842$\pm$ 0.0020.842 ± 0.002
1 Acquisition Sample 0.728±0.0010.728$\pm$ 0.0010.728 ± 0.001 0.699±0.0000.699$\pm$ 0.0000.699 ± 0.000 0.694±0.0030.694$\pm$ 0.0030.694 ± 0.003 0.827±0.0020.827$\pm$ 0.0020.827 ± 0.002
1 Train Sample 0.741±0.0020.741$\pm$ 0.0020.741 ± 0.002 0.707±0.0010.707$\pm$ 0.0010.707 ± 0.001 0.705±0.0030.705$\pm$ 0.0030.705 ± 0.003 0.834±0.0020.834$\pm$ 0.0020.834 ± 0.002
Deterministic Encoder 0.741±0.0020.741$\pm$ 0.0020.741 ± 0.002 0.704±0.0010.704$\pm$ 0.0010.704 ± 0.001 0.702±0.0040.702$\pm$ 0.0040.702 ± 0.004 0.838±0.0020.838$\pm$ 0.0020.838 ± 0.002
Feature Space Calculation 0.697±0.0010.697$\pm$ 0.0010.697 ± 0.001 0.634±0.0010.634$\pm$ 0.0010.634 ± 0.001 0.695±0.0010.695$\pm$ 0.0010.695 ± 0.001 0.795±0.0010.795$\pm$ 0.0010.795 ± 0.001
WO Copula $\underline{0.762\pm 0.001}$ 0.717±0.0010.717$\pm$ 0.0010.717 ± 0.001 $\underline{0.710\pm 0.002}$ 0.843±0.0020.843$\pm$ 0.0020.843 ± 0.002
WO Normalization $\underline{0.762\pm 0.001}$ 0.718±0.0010.718$\pm$ 0.0010.718 ± 0.001 0.709±0.0020.709$\pm$ 0.0020.709 ± 0.002 $\underline{0.844\pm 0.002}$
WO Prob Weighting 0.751±0.0010.751$\pm$ 0.0010.751 ± 0.001 0.694±0.0010.694$\pm$ 0.0010.694 ± 0.001 0.698±0.0030.698$\pm$ 0.0030.698 ± 0.003 0.832±0.0010.832$\pm$ 0.0010.832 ± 0.001
SEFA (full) 0.761±0.0010.761$\pm$ 0.0010.761 ± 0.001 0.721±0.0000.721$\pm$ 0.0000.721 ± 0.000 0.709±0.0030.709$\pm$ 0.0030.709 ± 0.003 0.843±0.0020.843$\pm$ 0.0020.843 ± 0.002

### v3

@@ 4 Limitations of CMI Maximization
First, greedy CMI maximization makes myopic acquisitions, which in some scenarios is guaranteed to be sub-optimal. We prove this with an example. Consider a feature vector with $d+1$ features, the first $d$ of which are binary, and the last taking an integer value from $1$ to $d$: $X\in\{0,1\}^{d}\times[d]$. The final feature acts as an indicator, informing us which of the other $d$ features gives the label, $y=x_{x_{d+1}}$. The optimal strategy is to first choose the indicator then its designated feature, requiring two acquisitions. However, the expected number of acquisitions to arrive at the same prediction by greedily maximizing CMI is $3-\frac{1}{d}$ (proof in Appendix F). The theoretical insight into why CMI fails is because possible future observations are not considered in the present decision since they are marginalized out, $p(x_{i},y|\mathbf{x}_{O})=\int p(x_{j},x_{i},y|\mathbf{x}_{O})dx_{j}$. Each acquisition is made as if there are no subsequent acquisitions and therefore the indicator is not chosen first. This is not specific to CMI, but any scoring that marginalizes out unobserved features.
@@ 5.1 Architecture and Training
Training.   The SEFA architecture is trained in a supervised manner, minimizing the predictive negative log-likelihood, avoiding potential issues associated with training using RL. We also impose an information bottleneck regularization term on the latent space $I(X_{S};Z)$ (Tishby et al., 1999), minimizing how much information about the features is encoded in the latent variable. The regularization term is implemented using a variational upper bound (Alemi et al., 2017), as the expected KL divergence between $p_{\theta}(\mathbf{z}|\mathbf{x}_{S})$ and a prior $p(\mathbf{z})$. We use $\mathcal{N}(0,1)$ as the prior so that the KL divergence has a closed-form expression. This is multiplied by $\beta$ to balance predictive performance with regularization. Additionally, to train SEFA to make predictions on arbitrary feature subsets, at each training step we first uniformly sample a removal probability, and then each feature has that probability of being removed. This gives the loss
$L=\mathop{{}{\mathbb{E}}}_{p_{\text{Data}}(\mathbf{x}_{S},y)}$ $\mathop{{}{\mathbb{E}}}_{p_{\text{Subsample}}(S^{\prime})}\bigg{[}-\log\big{(}p_{\theta,\phi}(y|\mathbf{x}_{S\cap S^{\prime}})\big{)}$ (1)
@@ 5.2 Acquisition
[…]
@@
In order to convert the gradients of $\mathbf{z}$ to feature scores, we calculate the length of the gradient in a feature’s associated latent components, $\mathbf{g}_{\mathcal{G}_{i}}||_{2}$. Note that this is only possible because we used feature-wise encoders rather than one fully connected encoder. Finally, we normalize scores to treat each latent sample equally, removing the effect of the total gradient length. See Figure 1 for an example calculation.
@@ 5.3 Understanding the Latent Space
Revisiting the Loss.   The predictive part of the loss in (1) is given by the negative log-likelihood $\log\mathop{{}{\mathbb{E}}}_{p_{\theta}(\mathbf{z}|\mathbf{x}_{S})}p_{\phi}(y|\mathbf{z})$. It is standard to take one latent sample to calculate this. However, we take multiple samples for two reasons. (1) The predictor network must see enough latent samples during training to learn to adapt based on different samples during acquisition. (2) If only one sample is taken for a given $\mathbf{x}_{S}$, the predictive loss encourages each sample to output the same prediction. Either the encoder $p_{\theta}(\mathbf{z}|\mathbf{x}_{S})$ reduces the latent standard deviation, or the predictor $p_{\phi}(y|\mathbf{z})$ predicts the same label distribution for the different latent samples. Either way, there is a lack of diversity in the latent space predictions. If multiple samples are taken, this allows individual samples’ predictions to vary more. This will not improve the predictions, but improves the acquisition since a more diverse set of possible realizations are used in the inner loop of the objective. We verify this empirically in Section 6 and Appendix C.
@@ 6 Experiments
ACFlow $7.730\pm 0.139$ $7.527\pm 0.254$ $9.194\pm 0.278$
DIME $4.079\pm 0.064$ $4.581\pm 0.217$ $5.667\pm 0.038$
EDDI $9.197\pm 0.203$ $9.214\pm 0.415$ $9.794\pm 0.186$
Fixed MLP $6.009\pm 0.000$ $5.996\pm 0.000$ $7.999\pm 0.000$
GDFS $568\pm 0.219 4.4$$484\pm 0.159$ $5.587\pm 0.201$
GSMRL $5.570\pm 0.127$ $6.227\pm 0.185$ $8.199\pm 0.067$
Opportunistic RL $4.201\pm 0.041$ $4.846\pm 0.021$ $5.850\pm 0.072$
Random $9.484\pm 0.006$ $499\pm 0.005 9.9$$987\pm 0.008$
VAE $589\pm 0.088 6.6$$667\pm 0.140$ $7.888\pm 0.065$
@@
$\beta=0$ $4.520\pm 0.082$ $4.578\pm 0.109$ $5.716\pm 0.104$
1 Acquisition Sample $4.683\pm 0.030$ $4.862\pm 0.035$ $5.700\pm 0.027$
1 Train Sample $4.421\pm 0.174$ $4.713\pm 0.160$ $5.188\pm 0.107$
Deterministic Encoder $4.593\pm 0.236$ $4.773\pm 0.228$ $5.744\pm 0.034$
Feature Space Calculation $5.111\pm 0.070$ $5.461\pm 0.128$ $5.977\pm 0.052$
WO Normalization $4.036\pm 0.009$ $4.104\pm 0.008$ $5.101\pm 0.015$
@@ 6.2 Datasets with Unknown Feature Orderings
ACFlow $0.899\pm 0.001$ $0.823\pm 0.013$ $0.558\pm 0.010$ $0.881\pm 0.018$
DIME $0.901\pm 0.001$ $0.907\pm 0.002$ $0.661\pm 0.002$ $0.951\pm 0.001$
EDDI $0.764\pm 0.005$ $0.705\pm 0.010$ $0.412\pm 0.013$ $0.842\pm 0.009$
Fixed MLP $0.883\pm 0.001$ $0.909\pm 0.001$ $0.658\pm 0.002$ $0.954\pm 0.000$
GDFS $0.900\pm 0.000$ $0.907\pm 0.001$ $0.653\pm 0.002$ $0.949\pm 0.000$
GSMRL $0.891\pm 0.001$ $0.879\pm 0.006$ $0.638\pm 0.003$ $0.946\pm 0.001$
Opportunistic RL $0.901\pm 0.000$ $0.910\pm 0.000$ $0.657\pm 0.001$ $0.953\pm 0.000$
Random $0.699\pm 0.001$ $0.816\pm 0.003$ $0.569\pm 0.003$ $0.912\pm 0.001$
VAE $0.901\pm 0.001$ $0.878\pm 0.002$ $0.633\pm 0.005$ $0.925\pm 0.002$
@@
ACFlow $0.667\pm 0.003$ $0.652\pm 0.002$ $0.542\pm 0.006$ $0.711\pm 0.009$
DIME $0.731\pm 0.002$ $0.703\pm 0.002$ $0.670\pm 0.007$ $0.805\pm 0.002$
EDDI $0.572\pm 0.003$ $0.604\pm 0.001$ $0.563\pm 0.011$ $0.634\pm 0.005$
Fixed MLP $0.708\pm 0.001$ $0.690\pm 0.001$ $0.685\pm 0.003$ $0.799\pm 0.004$
GDFS $0.732\pm 0.001$ $0.692\pm 0.002$ $0.671\pm 0.005$ $0.797\pm 0.002$
GSMRL $0.701\pm 0.002$ $0.683\pm 0.001$ $0.665\pm 0.002$ $0.781\pm 0.003$
Opportunistic RL $0.740\pm 0.000$ $0.708\pm 0.000$ $0.706\pm 0.004$ $0.838\pm 0.002$
Random $0.661\pm 0.001$ $0.648\pm 0.001$ $0.647\pm 0.005$ $0.753\pm 0.003$
VAE $0.716\pm 0.001$ $0.685\pm 0.001$ $0.690\pm 0.004$ $0.800\pm 0.003$
@@ Acknowledgements
The results in the TCGA experiment are based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga. We thank the ICML reviewers for their comments and suggestions. We also thank Lucie Charlotte Magister and Iulia Duta for providing feedback. Alexander Norcliffe is supported by a GSK plc grant. Changhee Lee is supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00358602) and the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT), Artificial Intelligence Graduate School Program (No. RS-2019-II190079, Korea University) and the Artificial Intelligence Star Fellowship Support Program to nurture the best talents (No. RS-2025-02304828).
@@ Appendix C Real Data Ablations
$\beta=0$ $\underline{0.906\pm 0.000}$ $0.917\pm 0.001$ $\underline{0.680\pm 0.005}$ $0.957\pm 0.000$
1 Acquisition Sample $0.889\pm 0.001$ $0.912\pm 0.001$ $0.667\pm 0.004$ $0.952\pm 0.000$
1 Train Sample $0.901\pm 0.002$ $0.912\pm 0.003$ $0.676\pm 0.006$ $0.953\pm 0.000$
Deterministic Encoder $0.900\pm 0.000$ $0.893\pm 0.006$ $0.667\pm 0.005$ $0.944\pm 0.002$
Feature Space Calculation $97\pm 0.002 0.880$$80\pm 0.003$ $0.600\pm 0.004$ $0.914\pm 0.004$
WO Copula $\underline{0.905\pm 0.000}$ $0.917\pm 0.002$ $0.675\pm 0.002$ $0.952\pm 0.001$
WO Normalization $\underline{0.905\pm 0.000}$ $0.919\pm 0.001$ $0.676\pm 0.005$ $0.957\pm 0.000$
WO Prob Weighting $0.904\pm 0.001$ $0.919\pm 0.001$ $0.674\pm 0.005$ $0.957\pm 0.000$
SEFA (full) $0.904\pm 0.001$ $0.919\pm 0.001$ $0.676\pm 0.005$ $0.957\pm 0.000$
@@
$\beta=0$ $0.759\pm 0.000$ $0.719\pm 0.000$ $0.706\pm 0.003$ $0.842\pm 0.002$
1 Acquisition Sample $0.728\pm 0.001$ $0.699\pm 0.000$ $0.694\pm 0.003$ $0.827\pm 0.002$
1 Train Sample $41\pm 0.002 0.707$$07\pm 0.001$ $0.705\pm 0.003$ $0.834\pm 0.002$
Deterministic Encoder $0.741\pm 0.002$ $0.704\pm 0.001$ $0.702\pm 0.004$ $0.838\pm 0.002$
Feature Space Calculation $0.697\pm 0.001$ $0.634\pm 0.001$ $0.695\pm 0.001$ $0.795\pm 0.001$
WO Copula $\underline{0.762\pm 0.001}$ $0.717\pm 0.001$ $\underline{0.710\pm 0.002}$ $0.843\pm 0.002$
WO Normalization $\underline{0.762\pm 0.001}$ $0.718\pm 0.001$ $0.709\pm 0.002$ $\underline{0.844\pm 0.002}$
WO Prob Weighting $0.751\pm 0.001$ $0.694\pm 0.001$ $0.698\pm 0.003$ $0.832\pm 0.001$
SEFA (full) $0.761\pm 0.001$ $0.721\pm 0.000$ $0.709\pm 0.003$ $0.843\pm 0.002$